How to find uniqueness of solution using characteristics curve and initial condition


1
Expert's answer
2021-10-12T10:23:19-0400

Solution:

Let's take the problem from the system of characteristic ODEs that :

"\frac{d x}{p}=\frac{d y}{1}=\frac{d u}{u}"

A first characteristic equation comes from "\frac{d x}{p}=\frac{d y}{1}" :

"x-p y=c_{1}"

A second characteristic equation comes from "\quad \frac{d y}{1}=\frac{d u}{u}" :

"u e^{-y}=c_{2}"

The general solution of the PDE expressed on the form of an implicit equation is:

"\Phi\left(x-p y, u e^{-y}\right)=0"

"\Phi" is an arbitrary function of two variables.

Equivalently, on explicit form :

"u e^{-y}=F(x-p y)"

F is an arbitrary function.

So, the general solution is :

"u(x, y)=e^{y} F(x-p y)"


Boundary condition:

"u(X, 0)=g(X)=e^{0} F(X-0)=F(X) \quad" any X, doesn't matter the notation of the variable. Now the function F(X) is known :

F(X)=g(X)

We put it into the above general solution where X=x-p y.

The particular solution which satisfies the boundary condition is :

"u(x, y)=e^{y} g(x-p y)"